Circular No. 8281

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

S/2003 J 23 Further to IAUC 8276, the discovery of a new satellite of Jupiter has

been reported on CCD images taken with three telescopes at Mauna Kea by a team led by S. S. Sheppard. Astrometry, preliminary orbital elements by B. G. Marsden, and an ephemeris appear on MPEC 2004-B81 (e = 0.31, i = 149^{o}, P = 759 days, H = 16.7).

SUPERNOVA 2004Q IN ESO 507-G11 Further to IAUC 8277, J. Burket and W. Li report the LOSS discovery,

on KAIT images taken on Jan. 30.5 and 31.5 UT, of an apparent supernova (mag ∼ 18.4) located at α = 12^{h}47^{m}40^{s}.11, δ = −26^{o}12′23′′.3 (equinox 2000.0), which is 2′′.5 west and 24′′.1 south of the nucleus of ESO 507-G11. A KAIT image taken on 2003 June 6.2 showed nothing at this position (limiting mag ∼ 19.0).

SUPERNOVAE 2003lt AND 2004R L. Strolger and A. Riess, Space Telescope Science Institute, report the

discovery of two apparent supernovae found in the observations of the Ultra Deep Field (UDF) and UDF Parallel fields on CCD images taken with the Advanced Camera for Surveys (ACS) on the Hubble Space Telescope. SN 2004R, which was detected by subtracting groups of z-band (F850LP) ACS images of the same fields, is located at α = 3^{h}32^{m}31^{s}.30, δ = −27^{o}46′13′′.6 (equinox 2000.0), which is 0′′.13 west and 0′′.10 south of the center of the apparent host galaxy. This new object was found in sixteen independent detection images taken during Jan. 2–11 and was not seen in images taken during 2003 Dec. 24–30 (limiting mag 28.5). The morphology of the transient is consistent with the ACS point-spread function (FWHM 0′′.1) and the lack of a detectable motion (< 0′′.05) relative to other sources over several hours of observation. Available z-band magnitude estimates: Jan. 10.44 UT, 27.2; 13.04, 26.5.

SN 2003lt was detected in eight separate exposures obtained during 2003 Aug. 30–Sept. 1, and it is located at α = 3^{h}32^{m}42^{s}.81, δ = −27^{o}55′53′′.1 (equinox 2000.0), which is 0′′.13 east and 0′′.50 north of the center of its apparent host galaxy. Available z-band magnitudes for SN 2003lt: 2003 Feb. 10.12 UT, [28.5, Sept. 1.89, 26.0; 13.62, 26.2.

2004 February 4  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2004 CBAT  Daniel W. E. Green